# Georgia Tech

### v1

@@ Abstract
Multimodal Large Language Models (MLLMs) hold significant promise for revolutionizing traditional education and reducing teachers’ workload. However, accurately interpreting unconstrained STEM student handwritten solutions with intertwined mathematical formulas, diagrams, and textual reasoning poses a significant challenge due to the lack of authentic and domain-specific benchmarks. Additionally, current evaluation paradigms predominantly rely on the outcomes of downstream tasks (e.g., auto-grading), which often probe only a subset of the recognized content, thereby failing to capture the MLLMs’ understanding of complex handwritten logic as a whole. To bridge this gap, we release EDU-CIRCUIT-HW, a dataset consisting of 1,300+ authentic student handwritten solutions from a university-level STEM course. Utilizing the expert-verified verbatim transcriptions and grading reports of student solutions, we simultaneously evaluate various MLLMs’ upstream recognition fidelity and downstream auto-grading performance. Our evaluation uncovers an astonishing scale of latent failures within MLLM-recognized student handwritten content, highlighting the models’ insufficient reliability for auto-grading and other understanding-oriented applications in high-stakes educational settings. In solution, we present a case study demonstrating that leveraging identified error patterns to preemptively detect and rectify recognition errors, with only minimal human intervention (approximately 4% of the total solutions), can significantly enhance the robustness of the deployed AI-enabled grading system on unseen student solutions.
@@ C Calculation Error: Wrong arithmetic with the correct equation
Figure 3: The proposed auto-grading pipeline. The “vanilla grading pipeline” in green box is a widely-used auto-grading paradigm in AI-enabled education field. In our experiment (Section 4), we detect and analyze recognition errors within “MLLM Recognized Text” utilizing the automated pipeline introduced in Section 3.1, and further evaluate “Grading Report” by benchmarking it against the “Expert’s Grading Report” in yellow box. In pink box, we present an exemplary regrading module where we summarize the detected recognition error patterns and leverage it to filter out potential recognition errors within “MLLM Recognized Text”, and let LLM/human regraders to reassess those solutions flagged with containing potential recognition error, more details can be found in Section 5.
@@ D.2 Vanilla Grading Pipeline
The “vanilla grading pipeline”, illustrated in Figure 3, comprises two main stages: (1) transcription of the student handwritten solution by an MLLM, such as Gemini-2.5-Pro, and (2) automated evaluation by an LLM-based grader. In the transcription stage, we utilize different MLLMs to convert handwritten student responses into structured text, guided by prompts specific to each model series (see Figures 14 and 15).
@@ D.3 Regrading Module
To implement this module, we first aggregated all recognition errors flagged by human experts within the observation set; the process for collecting these errors is detailed in Section C. Leveraging this prior knowledge, experts distilled a set of empirical criteria, presented in Figures 23 and 24, to guide the detection of potential recognition errors in the test set. These criteria were then integrated into the recognition error detector, which identifies likely transcription failures in student responses that received score deductions in the initial grading stage. For the case study, we use GPT-5.1 as the detector. Based on its output, the LLM regrader re-evaluates the affected student work using the prompt shown in Figure 22. During this process, the regrader is explicitly instructed to disregard any content flagged as a recognition error, even if it introduces discrepancies with the reference solution.

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) hold significant promise for revolutionizing traditional education and reducing teachers’ workload. However, accurately interpreting unconstrained STEM student handwritten solutions with intertwined mathematical formulas, diagrams, and textual reasoning poses a significant challenge due to the lack of authentic and domain-specific benchmarks. Additionally, current evaluation paradigms predominantly rely on the outcomes of downstream tasks (e.g., auto-grading), which often probe only a subset of the recognized content, thereby failing to capture the MLLMs’ understanding of complex handwritten logic as a whole. To bridge this gap, we release EDU-CIRCUIT-HW, a dataset consisting of 1,300+ authentic student handwritten solutions from a university-level STEM course. Utilizing the expert-verified verbatim transcriptions and grading reports of student solutions, we simultaneously evaluate various MLLMs’ upstream recognition fidelity and downstream auto-grading performance. Our evaluation uncovers an astonishing scale of latent failures within MLLM-recognized student handwritten content, highlighting the models’ insufficient reliability for auto-grading and other understanding-oriented applications in high-stakes educational settings. In solution, we present a case study demonstrating that leveraging identified error patterns to preemptively detect and rectify recognition errors, with only minimal human intervention (e.g., with 3.3% assignments routed to human graders while the rest to GPT-5.1 grader), can effectively enhance the robustness of the deployed AI-enabled grading system on unseen student solutions.
@@ C Calculation Error: Wrong arithmetic with the correct equation
Figure 3: The proposed auto-grading pipeline. The “vanilla grading pipeline” in green box is a widely-used auto-grading paradigm in AI-enabled education field. In our experiment (Section 4), we detect and analyze recognition errors within “MLLM Recognized Text” utilizing the automated pipeline introduced in Section 3.1, and further evaluate “Grading Report” by benchmarking it against the “Expert’s Grading Report” in yellow box. In pink box, we present an exemplary regrading module where we let the human experts summarize the detected recognition error patterns and leverage it ((detailed in Figure 25 and 26)) to filter out potential recognition errors within “MLLM Recognized Text”, and let LLM/human regraders to reassess those solutions flagged with containing potential recognition error, more details can be found in Section 5.
@@ 6 Conclusions
7 Data Scope: Beyond Circuit Analysis
As introduced in Section 2.1, the student handwritten data in EDU-CIRCUIT-HW was collected from a circuit-analysis-related university-level course. However, we argue that this does not imply that our data merely serves as a proxy for a narrow subset of STEM challenges. In effect, solving the problems included in our dataset requires interleaved reasoning across multiple foundational disciplines, including calculus, differential equations, and physics. These subjects are formal prerequisites for the course at our institution, meaning that students must demonstrate competence in them before enrolling in the target course. In Figure 10 and 11, we provide two recognized handwritten solutions corresponding to Problems 9.8-2 and 10.6-4 in the textbook by Svoboda and Dorf Svoboda and Dorf (2013). These two examples indicate that our EDU-CIRCUIT-HW dataset is also relevant to foundational STEM subjects such as physics, differential equations, calculus, and linear algebra. Due to copyright restrictions, we do not provide the original problem statements; instead, we briefly summarize them as follows:
i) The first example is a recognized handwritten solution to Problem 9.8-2 in the textbook [1]. The circuit diagram includes two resistors, one capacitor, one inductor, and an independent voltage source with voltage 2u(t)12u(t)-1. The task is to determine the current $i(t)$ through the inductor for $t>0$. The student’s solution involves fundamental circuit laws, including Kirchhoff’s Current Law (KCL) and Kirchhoff’s Voltage Law (KVL), the formulation and solution of a second-order linear differential equation, derivation of the characteristic equation in the Laplace domain, and representation of the final current as the sum of two exponential terms and one constant term. These elements indicate that the benchmarking results obtained from our dataset are not limited to circuit analysis content alone.
ii) The second example is a recognized handwritten solution to Problem 10.6-4 in the textbook [1]. The circuit diagram contains an independent voltage source represented by the phasor $48\angle 75^{\circ}$ and five impedance elements with impedances $40+j15\Omega$, $25-j50\Omega$, $32+j16\Omega$, $j50\Omega$, and $j40\Omega$. The task is to determine the values of three phasor mesh currents. The student’s solution involves three mesh-current equations forming a system of linear equations. When expressed in matrix form, the system includes complex-valued coefficients. Recognition and evaluation of this solution therefore require not only circuit analysis knowledge but also competence in linear algebra.
We hope the above clarification of our data scope can help the following researchers better understand and utilize our EDU-CIRCUIT-HW dataset.
@@ D.2 Vanilla Grading Pipeline
The “vanilla grading pipeline”, illustrated in Figure 3, comprises two main stages: (1) transcription of the student handwritten solution by an MLLM, such as Gemini-2.5-Pro, and (2) automated evaluation by an LLM-based grader. In the transcription stage, we utilize different MLLMs to convert handwritten student responses into structured text, guided by prompts specific to each model series (see Figures 16 and 17).
@@ D.3 Regrading Module
To implement this module, we first aggregated all recognition errors flagged by human experts within the observation set; the process for collecting these errors is detailed in Section C. Leveraging this prior knowledge, experts distilled a set of empirical criteria, presented in Figures 25 and 26, to guide the detection of potential recognition errors in the test set. These criteria were then integrated into the recognition error detector, which identifies likely transcription failures in student responses that received score deductions in the initial grading stage. For the case study, we use GPT-5.1 as the detector. Based on its output, the LLM regrader re-evaluates the affected student work using the prompt shown in Figure 24. During this process, the regrader is explicitly instructed to disregard any content flagged as a recognition error, even if it introduces discrepancies with the reference solution.
@@ D.5 Examples about Recognition Error Taxonomy
Example 1 for data scope clarification (A recognized handwritten solution to Problem 9.8-2; involving foundations from physics, differential equations, and calculus) For $t<0$: $i(0)=\frac{-1}{1+4}=-0.2\text{ A}$ $v(0)=\frac{4}{1+4}(-1)=-0.8\text{ V}$ For $t>0$: KCL $\rightarrow$ $\frac{v-V_{s}}{R_{1}}+C\frac{dv}{dt}+i=0$ KVL to right half $\rightarrow$ $v=R_{2}i+L\frac{di}{dt}$ plug KVL eq. to KCL eq.: $\frac{R_{2}i+L\frac{di}{dt}-V_{s}}{R_{1}}+C\left[R_{2}\frac{di}{dt}+L\frac{d^{2}i}{dt^{2}}\right]+i=0$ $\Rightarrow\frac{d^{2}i}{dt^{2}}+\frac{(L+R_{1}R_{2}C)}{R_{1}LC}\frac{di}{dt}+\frac{(R_{1}+R_{2})}{R_{1}LC}i=\frac{V_{s}}{R_{1}LC}$ $\Rightarrow\frac{d^{2}i}{dt^{2}}+5\frac{di}{dt}+5i=1$ Forced response is constant, $i_{f}=B$ so $1=\frac{d^{2}B}{dt^{2}}+5\frac{dB}{dt}+5B$ Therefore $B=0.2\text{ A}$ Characteristic eq: $s^{2}+5s+5=0$ Roots $\Rightarrow s_{1}=-3.62$, $s_{2}=-1.38$ $i_{n}=A_{1}e^{-3.62t}+A_{2}e^{-1.38t}$ $i(t)=A_{1}e^{-3.62t}+A_{2}e^{-1.38t}+0.2$ $v(t)=\left(4i(t)+4\frac{di(t)}{dt}\right)=-10.48A_{1}e^{-3.62t}-1.52A_{2}e^{-1.38t}+0.8$ At $t=0^{+}$, $0.2=i(0^{+})=A_{1}+A_{2}+0.2\Rightarrow A_{1}+A_{2}=-0.4$ $0.8=v(0^{+})=-10.48A_{1}-1.52A_{2}+0.8\Rightarrow-10.48A_{1}-1.52A_{2}=-1.6$ Using matrix calc: $A_{1}=0.246$ and $A_{2}=-0.646$ $i(t)=0.246e^{-3.62t}-0.646e^{-1.38t}+0.2\text{ A}$
Figure 10: An example in our dataset including a student’s solution on using foundations from physics, differential equations, and calculus to solve the target current.
Example 2 for data scope clarification (A recognized handwritten solution to Problem 10.6-4) Mesh 1: $(40+j15)i_{1}+(25-j50)(i_{1}-i_{3})-48\angle 75^{\circ}=0$ $(65-j35)i_{1}-(25-j50)i_{3}=48\angle 75^{\circ}$ Mesh 2: $48\angle 75^{\circ}+(-j50)(i_{2}-i_{3})+(32+j16)i_{2}=0$ $(32-j34)i_{2}+j50i_{3}=-48\angle 75^{\circ}$ Mesh 3: $j40i_{3}-(-j50)(i_{2}-i_{3})-(25-j50)(i_{1}-i_{3})=0$ $(-25+j50)i_{1}+j50i_{2}+(25-j160)i_{3}=0$ $\begin{bmatrix}-25+j50&j50&25-j160\\ 0&32-j34&j50\\ 65-j35&0&-25+j50\end{bmatrix}\begin{bmatrix}i_{1}\\ i_{2}\\ i_{3}\end{bmatrix}=\begin{bmatrix}0\\ -48\angle 75^{\circ}\\ 48\angle 75^{\circ}\end{bmatrix}$ $i_{1}=0.794\angle 111^{\circ}\text{ A}$,  $i_{2}=0.790\angle-61.7^{\circ}\text{ A}$,  $i_{3}=0.229\angle 176^{\circ}\text{ A}$
Figure 11: An example in our dataset including a student’s solution on using linear algebra and complex operation to solve the target currents.